# Forbidden Electron Transfer in the Adiabatic Limit of the Marcus-Inverted Region

### v1

@@ Abstract
To show (iii), which may be the most remarkable and counterintuitive proposition of this Letter, we note that it is not guaranteed that when the coupling is turned, that the electron will end up on the lower energy adiabat, even in the strong coupling limit. If the reactant coordinate is not near the diabatic crossing when the coupling is turned on, it will remain on the adiabat that most closely resembles its initial diabat regardless of whether it is the higher or lower energy one. For the same reason that ET in Marcus theory is assumed to be an isoenergetic process satisfying the Franck Condon principle [7], the system will remain trapped on the higher diabat and no reaction will occur, with the exception of rare nuclear tunneling events [19, 10]. In this system, the criteria for the higher-energy adiabat coinciding with the $D$ diabat is precisely the criteria of being in the inverted region of the diabatic Marcus picture, because this is the point at which the mapping between the diabats and adiabats is altered.

### v2

@@ Abstract
To show (iii), which may be the most remarkable and counterintuitive proposition of this Letter, we note that it is not guaranteed that when the coupling is turned, that the electron will end up on the lower energy adiabat, even in the strong coupling limit. If the reactant coordinate is not near the diabatic crossing when the coupling is turned on, it will remain on the adiabat that most closely resembles its initial diabat regardless of whether it is the higher or lower energy one. For the same reason that ET in Marcus theory is assumed to be an isoenergetic process satisfying the Franck Condon principle [7], the system will remain trapped on the higher adiabat and no reaction will occur, with the exception of rare nuclear tunneling events [19, 10]. In this system, the criteria for the higher-energy adiabat coinciding with the $D$ diabat is precisely the criteria of being in the inverted region of the diabatic Marcus picture, because this is the point at which the mapping between the diabats and adiabats is altered.